# Beyond Empathy: Integrating Diagnostic and Therapeutic Reasoning with Large Language Models for Mental Health Counseling

### v1

@@ Abstract
Large language models (LLMs) hold significant potential for mental health support, capable of generating empathetic responses and simulating therapeutic conversations. However, existing LLM-based approaches often lack the clinical grounding necessary for real-world psychological counseling, particularly in explicit diagnostic reasoning aligned with standards like the DSM/ICD and incorporating diverse therapeutic modalities beyond basic empathy or single strategies. To address these critical limitations, we propose PsyLLM, the first large language model designed to systematically integrate both diagnostic and therapeutic reasoning for mental health counseling. To develop the PsyLLM, we propose a novel automated data synthesis pipeline. This pipeline processes real-world mental health posts, generates multi-turn dialogue structures, and leverages LLMs guided by international diagnostic standards (e.g., DSM/ICD) and multiple therapeutic frameworks (e.g., CBT, ACT, psychodynamic) to simulate detailed clinical reasoning processes. Rigorous multi-dimensional filtering ensures the generation of high-quality, clinically aligned dialogue data. In addition, we introduce a new benchmark and evaluation protocol, assessing counseling quality across four key dimensions: comprehensiveness, professionalism, authenticity, and safety. Our experiments demonstrate that PsyLLM significantly outperforms state-of-the-art baseline models on this benchmark.
@@
He Hu1,2111Equal Contribution., Yucheng Zhou3111Equal Contribution., Juzheng Si4, Qianning Wang5, Hengheng Zhang2 Fuji Ren6, Fei Ma2222Corresponding Authors., Laizhong Cui1,2222Corresponding Authors. 1College of Computer Science and Software Engineering, Shenzhen University 2Guangdong Laboratory of Artificial Intelligence and Digital Economy (SZ) 3SKL-IOTSC, CIS, University of Macau, 4Shandong University 5Auckland University of Technology, 6University of Electronic Science and Technology of China huhe@gml.ac.cn, yucheng.zhou@connect.um.edu.mo, mafei@gml.ac.cn, cuilz@szu.edu.cn
@@ 1 Introduction
| Dataset | Turn Type | Reasoning? | Therapy? | Diagnostic? | Dialogues |
| SweetieChat Ye et al. (2025) | Multi-turn | ✗ | ✗ | ✗ | 3,757 |
| ESD Zhang et al. (2024b) | Multi-turn | ✗ | ✗ | ✗ | 2,900 |
| CPsyCounD Zhang et al. (2024a) | Multi-turn | ✗ | ✓ | ✗ | 3,134 |
| ChatCounselor Liu et al. (2023a) | Single-turn | ✗ | ✗ | ✗ | 8,187 |
| OpenR1-Psy (Ours) | Single + Multi-turn | ✓ | ✓ | ✓ | 19,302 |
@@
To address existing limitations, we propose the first large language model that systematically integrates both diagnostic and therapeutic reasoning for mental health counseling, as shown in Fig. 1. Specifically, the model is capable of autonomously applying diagnostic criteria and various psychotherapeutic frameworks during counseling interactions. As shown in Table 1, we introduce a novel data synthesis pipeline that: (1) automatically parses real-world mental health posts; (2) generates simplified multi-turn diagnostic dialogues; (3) guides LLMs to simulate the reasoning processes of mental health professionals, incorporating step-by-step thinking aligned with DSM/ICD diagnostic logic and therapeutic considerations drawn from multiple evidence-based treatments; and (4) applies multi-dimensional filtering—including Incomplete Thinking, Context Incoherence, Mismatch Response, and No Framework—to ensure data quality. The resulting high-quality filtered dataset is used to fine-tune the LLM, enabling it to deliver counseling responses that are not only accurate and therapeutically grounded but also contextually appropriate and aligned with clinical reasoning processes.
In our experiments, we construct a benchmark to evaluate the quality of counseling responses along four key dimensions: comprehensiveness, professionalism, authenticity, and safety. Furthermore, we analyze the training data used for counseling responses, revealing that while increased model scale improves performance, the importance of data quantity is highly dependent on its quality. Our main contributions are as follows:
@@
We incorporate international diagnostic standards (e.g., DSM/ICD) and diverse therapeutic strategies (e.g., CBT, ACT, psychodynamic) to guide our model’s deep reasoning.
@@ 2.1 Mental Support Chatbots
The increasing demand for accessible mental health services has spurred the development of AI-driven dialogue systems offering emotional support Ge et al. (2025). Early efforts, such as ChatCounselor Liu et al. (2023b) and PsyQA Sun et al. (2021), focused on single-turn, empathy-oriented responses, lacking therapeutic modeling or conversational depth. Subsequent work addressed multi-turn interaction: SMILE Qiu et al. (2024a) used ChatGPT to expand single replies into coherent dialogues, while SoulChat Chen et al. (2023) fine-tuned LLMs for better emotional adaptation across turns. SweetieChat Ye et al. (2025) introduced strategy-guided role-play to simulate diverse emotional scenarios, and CPsyCoun Zhang et al. (2024a) annotated dialogues with therapeutic types (e.g., CBT, humanistic), though it did not model underlying reasoning. To improve interpretability, ESCoT Zhang et al. (2024b) proposed a strategy-driven chain-of-thought process—emotion recognition, appraisal, and justification—mirroring human counseling. MentaLLaMA Yang et al. (2024) combined psychological insight with clinical condition prediction and rationales in social media settings. At the clinical edge, PsyGUARD Qiu et al. (2024b) addressed suicide risk detection and safety-aware response generation in multi-turn counseling.
@@ 2.2 Deep Reasoning in LLMs
Chain-of-Thought (CoT) prompting Wei et al. (2022) revealed LLMs’ capacity for multi-step inference, significantly improving symbolic reasoning. This has since extended to decision-making, multimodal understanding, and emotional inference. Dual-system models Dou et al. (2025) and modality-aware reinforcement learning Zhao et al. (2025) enhance adaptability and emotion recognition. In math reasoning, Wang et al. (2025) aligns vision and code for structured geometry tasks, outperforming GPT-4o. Dialogue reasoning introduces challenges in coherence and emotional depth; Shu et al. (2025) reframes CoT as agent dialogue, while Xie et al. (2025) fuses CoT with emotional cues from speech. Psychological reasoning is supported by CoT-modeled questionnaires Yang et al. (2023) and graph-based therapist–AI collaboration tools Chen et al. (2025). The full version of the related work can be found in Appendix A.
@@ 3.1 Data Collection
To ensure both clinical realism and contextual diversity, our dataset sources are primarily divided into two parts: (1) The Reddit posts we used were pre-filtered and sourced from three publicly available corpora in mental health-related communities: Identifying-depression Pirina and Çöltekin (2018), Dreaddit Turcan and McKeown (2019), and LRF Garg et al. (2023). (2) Authentic patient utterances are derived from two high-quality psychological counseling datasets, ChatCounselor Liu et al. (2023b) and CPsyCoun Zhang et al. (2024a), which are based on real-world clinical cases and authentic patient-doctor dialogues, and refined as multi-turn dialogues between patients and LLM counselors by ChatGPT OpenAI (2023).
@@ 3.5 Systematical Dialogue Validation
After the generation process, synthesized dialogue turns are constructed by combining patient utterances with their corresponding generated reasoning traces and counselor responses . To ensure the clinical relevance, logical consistency, and overall quality of this data for fine-tuning PsyLLM, we implement a rigorous validation process. This validation employs a multi-dimensional filtering mechanism applied to each generated turn based on and context . A sample is kept if it satisfies the following boolean condition :
where the boolean variables represent the evaluation along four key dimensions: , Complete Thinking Detection, which assesses the completeness and logical flow of the reasoning trace ; , Context Coherence Evaluation, which evaluates the consistency of with the preceding dialogue context and the user prompt ; , Mismatch Response Checking, which verifies the alignment between the reasoning trace and the counselor response ; and , Reasoning Framework Checking, which confirms whether adheres to the intended guidance from international diagnostic standards () and therapeutic strategies ().
Through this process, low-quality generated segments are filtered out, resulting in a high-quality dataset used for training. This systematic quality control is essential for building a reliable and clinically informed mental health counseling system.
@@ 3.6 Dataset Statistics and Analysis
| # Dialogues | 18,852 | 450 | 19302 |
| # Utterances | 47,723 | 1651 | 49374 |
| Avg. Turn per Dialogue | 2.53 | 3.67 | 2.56 |
@@
| Avg. Reasoning Utterance Length | 1592.33 | 1627.36 | 1593.5 |
We analyze the constructed OpenR1-Psy dataset, presenting its key statistics and characteristics in Table 2 and Fig. 3. Table 2 shows the dataset statistics, comprising 19,302 dialogues (49,374 utterances) with a train/test split. The dialogues are multi-turn (avg. 2.56 turns). Patient utterances average 222 tokens, counselor responses 519 tokens, and notably, the reasoning trace averages 1593 tokens, highlighting its detailed nature. Fig. 3 presents key dataset distributions. (Left) shows psychotherapy approaches, dominated by Integrative (54.50%), Humanistic (25.15%), and CBT (17.16%), while including other diverse modalities. (Middle) illustrates scene categories, covering a broad range of topics like Emotion & Stress, Family Relationship, and Social Relationship. (Right) depicts severity levels: Moderate (48%) and Severe (41%) are most frequent, alongside Mild (10%) and Critical (1%), representing varied intensity.
@@ 4.1 Experimental Settings
We employ a combination of automatic metrics and expert human review to rigorously assess the quality of the generated counselor responses and their accompanying reasoning traces. Our evaluation framework focuses on four key dimensions: comprehensiveness, assessing if the response addresses the core issues; professionalism, evaluating adherence to ethical guidelines and therapeutic principles; authenticity, ensuring the response feels natural and empathetic; and safety, verifying the absence of harmful or inappropriate content. These dimensions are applied to evaluate each generated single turn, which includes both the reasoning trace and the counselor utterance. We apply strict, reality-aligned criteria within these dimensions to ensure our evaluation reflects the standards expected in real-world therapeutic practice for individual responses.
| Method | Metrics | ||||
| Comprehensiveness | Professionalism | Authenticity | Safety | Normalized Avg | |
| ChatCounselor Liu et al. (2023b) | 0.740 | 1.830 | 1.430 | 0.990 | 0.574 |
| CPsyCounD Zhang et al. (2024a) | 0.950 | 1.950 | 1.840 | 0.980 | 0.639 |
| OpenR1-Psy | 1.63 (+71%) | 3.59 (+84%) | 2.66 (+44%) | 1.00 (+2%) | 0.900 (+41%) |
| Method | Metrics | ||||
| Comprehensiveness | Professionalism | Authenticity | Safety | Normalized Avg | |
| ChatCounselor Liu et al. (2023b) | 0.015 | 0.826 | 0.332 | 0.469 | 0.202 |
| CPsyCounX Zhang et al. (2024a) | 0.005 | 0.492 | 0.180 | 0.358 | 0.137 |
| DeepSeek-V3 DeepSeek-AI et al. (2024) | 0.880 | 2.155 | 1.057 | 0.830 | 0.540 |
| GPT-4o Hurst et al. (2024) | 0.833 | 2.218 | 1.038 | 0.903 | 0.552 |
| PsyLLM | 0.904 | 2.210 | 1.440 | 0.945 | 0.607 |
| Method | Human | ||||
| Comprehensiveness | Professionalism | Authenticity | Safety | Normalized Avg | |
| ChatCounselor Liu et al. (2023b) | 0.280 | 0.950 | 0.630 | 0.870 | 0.364 |
| CPsyCounX Zhang et al. (2024a) | 0.120 | 0.480 | 0.270 | 0.600 | 0.204 |
| DeepSeek-V3 DeepSeek-AI et al. (2024) | 0.880 | 1.880 | 1.250 | 0.810 | 0.507 |
| GPT-4o Hurst et al. (2024) | 0.850 | 2.170 | 1.270 | 0.920 | 0.541 |
| PsyLLM | 0.990 | 2.330 | 1.620 | 0.980 | 0.601 |
@@ 4.2 Synthetic Data Comparison
We conducted a synthetic data comparison to evaluate the quality of data generated by different methods. We randomly selected 150 patient cases from ChatCounselor Liu et al. (2023b) and 150 cases from CPsyCounD Zhang et al. (2024a), comparing their original counselor responses with those generated by our OpenR1-Psy method for the same inputs. Gemini-2.0-Flash Anil et al. (2023) evaluated these responses based on the four metrics: Comprehensiveness, Professionalism, Authenticity, and Safety (detailed prompts in Fig. 10).
Table 3 presents the evaluation results. OpenR1-Psy achieved substantial improvements over ChatCounselor, with gains of 120% in Comprehensiveness, 96% in Professionalism, 86% in Authenticity, and 1% in Safety, resulting in a 57% higher normalized average. Compared to CPsyCounD, OpenR1-Psy showed notable gains of 71% (Comprehensiveness), 84% (Professionalism), 44% (Authenticity), 2% (Safety), and 41% higher normalized average. These results confirm that the OpenR1-Psy generation method produces significantly higher quality data across key evaluation dimensions.

### v2

@@ Abstract
Large language models (LLMs) hold significant potential for mental health support, capable of generating empathetic responses and simulating therapeutic conversations. However, existing LLM-based approaches often lack the clinical grounding necessary for real-world psychological counseling, particularly in explicit diagnostic reasoning aligned with standards like the DSM/ICD and incorporating diverse therapeutic modalities beyond basic empathy or single strategies. To address these critical limitations, we propose PsyLLM, the first large language model designed to systematically integrate both diagnostic and therapeutic reasoning for mental health counseling. To develop PsyLLM, we design a novel automated data synthesis pipeline that processes real-world mental health posts collected from Reddit, where users frequently share psychological distress and seek community support. This pipeline processes real-world mental health posts, generates multi-turn dialogue structures, and leverages LLMs guided by international diagnostic standards (e.g., DSM/ICD) and multiple therapeutic frameworks (e.g., CBT, ACT, psychodynamic) to simulate detailed clinical reasoning processes. Rigorous multi-dimensional filtering ensures the generation of high-quality, clinically aligned dialogue data. In addition, we introduce a new benchmark and evaluation protocol, assessing counseling quality across four key dimensions. Our experiments demonstrate that PsyLLM significantly outperforms state-of-the-art baseline models on this benchmark. The model weights and dataset have been publicly released at https://github.com/Emo-gml/PsyLLM.
@@
He Hu1,2111Equal Contribution. Yucheng Zhou3111Equal Contribution. Juzheng Si4 Qianning Wang5 Hengheng Zhang2 Fuji Ren2 Fei Ma2222Corresponding Authors. Laizhong Cui1,2222Corresponding Authors. Qi Tian2 1College of Computer Science and Software Engineering, Shenzhen University 2Guangdong Laboratory of Artificial Intelligence and Digital Economy (SZ) 3SKL-IOTSC, CIS, University of Macau 4Shandong University 5Auckland University of Technology 6University of Electronic Science and Technology of China huhe@gml.ac.cn, yucheng.zhou@connect.um.edu.mo, mafei@gml.ac.cn, cuilz@szu.edu.cn
@@ 1 Introduction
| Dataset | Publicly Available | Turn Type | Reasoning ? | Therapy ? | Diagnostic ? | Dialogues |
| Psych8k Liu et al. (2023a) | ✓ | Single-turn | ✗ | ✗ | ✗ | 8,187 |
| CPsyCounD Zhang et al. (2024a) | ✓ | Multi-turn | ✗ | ✓ | ✗ | 3,134 |
| SweetieChat Ye et al. (2025) | ✗ | Multi-turn | ✗ | ✗ | ✗ | 3,757 |
| PsyDTCorpus Xie et al. (2025a) | ✗ | Multi-turn | ✗ | ✗ | ✗ | 5,000 |
| MDD-5k Yin et al. (2025) | ✗ | Multi-turn | ✗ | ✗ | ✓ | 5,000 |
| OpenR1-Psy (Ours) | ✓ | Single + Multi-turn | ✓ | ✓ | ✓ | 19,302 |
@@
To address existing limitations, we propose the first large language model that systematically integrates both diagnostic and therapeutic reasoning for mental health counseling, as shown in Fig. 1. Specifically, the model is capable of autonomously applying diagnostic criteria and various psychotherapeutic frameworks during counseling interactions. As shown in Table 1, we introduce a novel data synthesis pipeline that: (1) automatically parses real-world mental health posts; (2) generates simplified multi-turn diagnostic dialogues; (3) guides LLMs to simulate the reasoning processes of mental health professionals, incorporating step-by-step thinking aligned with DSM/ICD diagnostic logic and therapeutic considerations drawn from multiple evidence-based treatments; and (4) applies multi-dimensional filtering, including Incomplete Thinking, Context Incoherence, Mismatch Response, and No Framework, to ensure data quality. The resulting high-quality filtered dataset is used to fine-tune the LLM, enabling it to deliver counseling responses that are not only accurate and therapeutically grounded but also contextually appropriate and aligned with clinical reasoning processes.
In our experiments, we construct a benchmark to evaluate the quality of counseling responses along four key dimensions: Empathy & Insight, Support & Autonomy, Attunement & Presence, Safety & Boundaries. Furthermore, we analyze the training data used for counseling responses, revealing that while increased model scale improves performance, the importance of data quantity is highly dependent on its quality. Our main contributions are as follows:
@@
We incorporate international diagnostic standards (e.g., DSM/ICD) and diverse therapeutic strategies (e.g., CBT, ACT, psychodynamic) to guide model’s deep reasoning.
@@ 2.1 Mental Support Chatbots
The increasing demand for accessible mental health services has spurred the development of AI-driven dialogue systems offering emotional support Ge et al. (2025). Early efforts, such as ChatCounselor Liu et al. (2023b) and PsyQA Sun et al. (2021), focused on single-turn, empathy-oriented responses, lacking therapeutic modeling or conversational depth. Subsequent work addressed multi-turn interaction: SMILE Qiu et al. (2024a) used ChatGPT to expand single replies into coherent dialogues, while SoulChat Chen et al. (2023) fine-tuned LLMs for better emotional adaptation across turns. SweetieChat Ye et al. (2025) introduced strategy-guided role-play to simulate diverse emotional scenarios, and CPsyCoun Zhang et al. (2024a) annotated dialogues with therapeutic types (e.g., CBT, humanistic), though it did not model underlying reasoning. To improve interpretability, ESCoT Zhang et al. (2024b) proposed a strategy-driven chain-of-thought process, emotion recognition, appraisal, and justification, mirroring human counseling. MentaLLaMA Yang et al. (2024) combined psychological insight with clinical condition prediction and rationales in social media settings. At the clinical edge, PsyGUARD Qiu et al. (2024b) addressed suicide risk detection and safety-aware response generation in multi-turn counseling. Existing chatbot-based approaches largely focus on empathetic responses or surface-level support, lacking clinically grounded diagnostic and therapeutic reasoning. In contrast, PsyLLM explicitly integrates both to support clinically informed counseling.
@@ 2.2 Deep Reasoning in LLMs
Chain-of-Thought (CoT) prompting Wei et al. (2022) revealed LLMs’ capacity for multi-step inference, significantly improving symbolic reasoning. This has since extended to decision-making, multimodal understanding, and emotional inference. Dual-system models Dou et al. (2025) and modality-aware reinforcement learning Zhao et al. (2025) enhance adaptability and emotion recognition. In math reasoning, Wang et al. (2025) aligns vision and code for structured geometry tasks, outperforming GPT-4o. Dialogue reasoning introduces challenges in coherence and emotional depth; Shu et al. (2025) reframes CoT as agent dialogue, while Xie et al. (2025b) fuses CoT with emotional cues from speech. Psychological reasoning is supported by CoT-modeled questionnaires Yang et al. (2023) and graph-based therapist–AI collaboration tools Chen et al. (2025). The full version of the related work can be found in Appendix A.
@@ 3.1 Data Collection
To ensure both clinical realism and contextual diversity, our dataset sources are primarily divided into two parts: (1) The Reddit posts we used were pre-filtered and sourced from three publicly available corpora in mental health-related communities: Identifying-depression Pirina and Çöltekin (2018), Dreaddit Turcan and McKeown (2019), and LRF Garg et al. (2023). (2) Authentic patient utterances are derived from two high-quality psychological counseling datasets, i.e., ChatCounselor Liu et al. (2023b) and CPsyCoun Zhang et al. (2024a), which are based on real-world clinical cases and authentic patient-doctor dialogues, and refined as multi-turn dialogues between patients and LLM counselors by ChatGPT OpenAI (2023).
@@ 3.5 Systematical Dialogue Validation
After the generation process, synthesized dialogue turns are constructed by combining patient utterances with their corresponding generated reasoning traces and counselor responses . To ensure the clinical relevance, logical consistency, and overall quality of this data for fine-tuning PsyLLM, we implement a rigorous validation process. This validation employs a multi-dimensional filtering mechanism applied to each generated turn based on and context . A sample is kept if it satisfies the boolean condition :
where the boolean variables represent the evaluation along four key dimensions: , Complete Thinking Detection, which assesses the completeness and logical flow of the reasoning trace ; , Context Coherence Evaluation, which evaluates the consistency of with the preceding dialogue context and the user prompt ; , Mismatch Response Checking, which verifies the alignment between the reasoning trace and the counselor response ; and , Reasoning Framework Checking, which confirms whether adheres to the intended guidance from international diagnostic standards () and therapeutic strategies (). Through this process, low-quality generated segments are filtered out, resulting in a high-quality dataset used for training.
@@ 3.6 Dataset Statistics and Analysis
| # Dialogues | 18,852 | 450 | 19,302 |
| # Utterances | 47,723 | 1,651 | 49,374 |
| Avg. Multi-turns | 3.76 | 5.04 | 3.80 |
@@
| Avg. Reasoning Utterance Length | 1,592.33 | 1,627.36 | 1,593.5 |
| Method | Metrics | ||||
| Emp&In | Sup&Aut | Att&Pre | Saf&Bou | Normalized Avg | |
| Psych8k Liu et al. (2023b) | 0.740 | 1.830 | 1.430 | 0.990 | 0.574 |
| CPsyCounD Zhang et al. (2024a) | 0.950 | 1.950 | 1.840 | 0.980 | 0.639 |
| OpenR1-Psy | 1.630 (+71%) | 3.590 (+84%) | 2.660 (+44%) | 1.000 (+2%) | 0.900 (+41%) |
We analyze the constructed OpenR1-Psy dataset, presenting its key statistics and characteristics in Table 2 and Fig. 3. Table 2 shows the dataset statistics, comprising 19,302 dialogues (49,374 utterances) with a train/test split. Patient utterances average 222 tokens, counselor responses 519 tokens, and notably, the reasoning trace averages 1593 tokens, highlighting its detailed nature. Fig. 3 presents key dataset distributions. (Left) shows psychotherapy approaches, dominated by Integrative (54.50%), Humanistic (25.15%), and CBT (17.16%), while including other diverse modalities. (Middle) illustrates scene categories, covering a broad range of topics like Emotion & Stress, Family Relationship, and Social Relationship. (Right) depicts severity levels: Moderate (48%) and Severe (41%) are most frequent, alongside Mild (10%) and Critical (1%), representing varied intensity.
@@ 4.1 Experimental Settings
We employ a combination of automatic metrics and expert human review to rigorously assess the quality of the generated counselor responses and their accompanying reasoning traces. Guided by prior research Yang et al. (2024); Zhang et al. (2024a); Xie et al. (2025a); Zhao et al. (2024). Our evaluation framework focuses on four key dimensions: Empathy & Insight (EmpIn), assessing recognition of core emotions and deeper unmet needs; Support & Autonomy (SupAut), evaluating natural, respectful support that fosters agency; Attunement & Presence (AttPre), ensuring conversational fluency, emotional connection, and responsiveness; and Safety & Boundaries (SafBou), verifying ethical sensitivity, psychological safety, and avoidance of intrusive or harmful content. These dimensions are applied to evaluate each generated single turn, which includes both the reasoning trace and the counselor utterance. We apply reality-aligned criteria in these dimensions to ensure our evaluation reflects the standards expected in real-world therapeutic practice for individual responses.
| Evaluator | Method | Metrics | ||||
| Emp&In | Sup&Aut | Att&Pre | Saf&Bou | Normalized Avg | ||
| Automatic | ChatCounselor Liu et al. (2023b) | 0.015 | 0.826 | 0.332 | 0.469 | 0.202 |
| CPsyCounX Zhang et al. (2024a) | 0.005 | 0.492 | 0.180 | 0.358 | 0.137 | |
| MeChat Qiu et al. (2024a) | 0.018 | 0.856 | 0.414 | 0.800 | 0.291 | |
| PsyDTLLM Xie et al. (2025a) | 0.148 | 1.135 | 0.689 | 0.872 | 0.365 | |
| DeepSeek-V3 DeepSeek-AI et al. (2024) | 0.880 | 2.155 | 1.057 | 0.830 | 0.540 | |
| GPT-4o Hurst et al. (2024) | 0.833 | 2.218 | 1.038 | 0.903 | 0.552 | |
| PsyLLM | 0.904 | 2.210 | 1.440 | 0.945 | 0.607 | |
| Human | ChatCounselor Liu et al. (2023b) | 0.280 | 0.950 | 0.630 | 0.870 | 0.364 |
| CPsyCounX Zhang et al. (2024a) | 0.120 | 0.480 | 0.270 | 0.600 | 0.204 | |
| MeChat Qiu et al. (2024a) | 0.280 | 1.350 | 0.690 | 0.820 | 0.382 | |
| PsyDTLLM Xie et al. (2025a) | 0.320 | 1.230 | 0.740 | 0.900 | 0.404 | |
| DeepSeek-V3 DeepSeek-AI et al. (2024) | 0.880 | 1.880 | 1.250 | 0.810 | 0.507 | |
| GPT-4o Hurst et al. (2024) | 0.850 | 2.170 | 1.270 | 0.920 | 0.541 | |
| PsyLLM | 0.990 | 2.330 | 1.620 | 0.980 | 0.601 | |
@@ 4.2 Synthetic Data Comparison
We conducted a synthetic data comparison to evaluate the quality of data generated by different methods. We randomly selected 150 patient cases from Psych8k Liu et al. (2023b) and 150 cases from CPsyCounD Zhang et al. (2024a), comparing their original counselor responses with those generated by our OpenR1-Psy method for the same inputs. Gemini-2.0-Flash Anil et al. (2023) evaluated these responses based on our four evaluation metrics.
Table 3 presents the evaluation results. OpenR1-Psy achieved substantial improvements over ChatCounselor, with gains of 120% in Emp&In, 96% in Sup&Aut, 86% in Att&Pre, and 1% in Saf&Bou, resulting in a 57% higher normalized average. Compared to CPsyCounD, OpenR1-Psy showed notable gains of 71% (Emp&In), 84% (Sup&Aut), 44% (Att&Pre), 2% (Saf&Bou), and 41% higher normalized average. These confirm that OpenR1-Psy generation method produces significantly higher quality data across key evaluation dimensions.